OpenStudy (anonymous):

The first side of a triangle measures 5 in. less than the second side, the third side is 3 in. more than the first side, and the perimeter is 17 in. If s represents the length of the second side, which of the following represents the length of the third side? s - 5 s - 2 s + 3

OpenStudy (anonymous):

s-2

OpenStudy (savvy):

s-2 since first side = s-5 third side = (s-5) + 3 = s-2